# Towards an Agnostic Algorithm for Sampling Empirical Structure Models

### v1

@@ 3.2 Derived quantities
Figure 6: Pressure Profiles. Weighted distribution of pressure profiles derived from all successful density profiles. The inset shows a zoomed in view of the outermost region, from $r/R=1$ to $r/R=0.7$. Note that the zoom is not aspect-preserving. The colour scale shows the relative frequency over all distributions of a certain pressure value at a given radius. The coloured lines are as before in Figures 2(a) and 3(a).
@@
In this Section, we present quantities derived from the successful density profiles via the Theory of Figures. Figure 6 shows the distribution of pressure profiles with a logarithmic frequency scale. Figure 7 shows the distribution of the normalised moments of inertia and Figure 8 shows the distribution of the flattening ratios.
We find that the pressure profiles in Figure 6 agree well with previous studies and are more constrained for Uranus compared to Neptune. The pressure profiles were calculated from the density profiles under the assumption of hydrostatic equilibrium $\nabla p=\rho\nabla U$, where $U$ is the total (gravitational and rotational) potential of the planet.
The normalised moments of inertia depicted in Figure 7 are defined as $\lambda=I/(mR_{\text{mean}}^{2})$, where $I$ is the axial moment of inertia
$I=\int_{V}\rho r^{2}_{\perp}\mathrm{d}V$ (7)
and $r_{\perp}$ denotes the perpendicular radial component with respect to the given axis, usually the axis of rotation. $R_{\text{mean}}$ is equal to $l_{\text{max}}$, the outermost spheroid volumetric mean radius, see Section 2. We find that our average value for Uranus of 0.22595 agrees well with previous findings from Neuenschwander & Helled (2022), who found a range of 0.22594–0.22670. Our average value for Neptune of 0.23597 is lower than their finding of 0.23727–0.23900 for the standard model, though our more frequent values agree with the lower end of their range.
@@
To compare our findings, we calculate the equatorial to mean volumetric radii ratios based on the values from Seidelmann et al. (2007). These are $(25559\pm 4)/(25362\pm 7)$ for Uranus and $(24764\pm 15)/(24622\pm 19)$ for Neptune. The resulting flattening ratios are $r_{f}=1.0078\pm 0.0003$ and $r_{f}=1.0058\pm 0.0010$ for Uranus and Neptune, respectively. Our inferred average value of 1.0068 in Figure 8 for Uranus is lower compared to Seidelmann et al. (2007), while our average of 1.0063 for Neptune is within their uncertainties. Our Theory of Figures implementation is benchmarked against Wisdom & Hubbard (2016) which shows that our flatting ratios should be accurate up to a relative error of $10^{-3}-10^{-4}$. This is significant given the uncertainties from Seidelmann et al. (2007) for Uranus and can therefore explain this deviation. Furthermore, the uncertainties given by Seidelmann et al. (2007) are quite conservative, especially given the uncertain rotation rates of these planets (for example Helled et al., 2010).
@@ 3.3 Discontinuities
Figure 9: Weighted discontinuity ‘intensity’ for Uranus (9(a)) and Neptune (9(b)). The figures depicts the average height $\overline{\Delta\rho}$ divided by the average width $\overline{\Delta r}$ of discontinuities. For example, Uranus discontinuities with an average location of $r/R=0.5$ possess $\overline{\Delta\rho/\Delta r}\sim 0.5\Delta\ell^{-1}$ kg m-3. Profiles with no discontinuities also contribute to the averages with value zero.
In this Section, we present an analysis on discontinuities of successful density profiles. First, a ‘discontinuity criterion’ $c_{D}$ must be chosen in order to distinguish discontinuities from regular slopes. Unless otherwise specified, all results were obtained with $c_{D}=$ 100 $\Delta\ell^{-1}$ kg m-3. As a reminder: $\Delta\ell$ is equal to 24960 m for Uranus and 24186 m for Neptune (rounded to the nearest meter). Therefore, $c_{D}$ has a value of 4.01 $\cdot$ 10-3 kg m-4 for Uranus and 4.13 $\cdot$ 10-5 kg m-4 for Neptune (rounded to two significant figures). As a second distinguishing feature, we only consider discontinuities that are ‘minimally maximal’. A precise definition of the term minimally maximal is given in Appendix D. In a nutshell, minimally maximal specifies that the (radial) width of a discontinuity can neither be extended (because it would add sections of the density profile below $c_{D}$) nor shortened (because it would remove sections of the density profile exceeding $c_{D}$).
Figure 9 shows the average discontinuity ‘intensity’ as a function of radius for all successful density profiles for both Uranus and Neptune. More precisely, it shows the average discontinuity height $\overline{\Delta\rho}$ divided by the average discontinuity width $\overline{\Delta r}$ for discontinuities that are minimally maximal and in agreement with $c_{D}$ at their respective average normalised radius $r/R$. For example, Figure 9(a) shows that minimally maximal and $c_{D}$-consistent discontinuities in Uranus with an average location of $r/R=0.5$ possess an average slope of $\overline{\Delta\rho/\Delta r}\sim 0.5\Delta\ell^{-1}$ kg m-3. Note that profiles that have no minimally maximal and $c_{D}$-consistent discontinuity at any given location also contribute to the averages with discontinuity intensity zero. We separately show $\overline{\Delta\rho}$ and $\overline{\Delta r}$ at their respective average normalised radii $r/R$ in Figures 14 for Uranus and 15 for Neptune.
@@
For Uranus (Figure 9(a)), we observe a discontinuity intensity peak at $\sim 0.65R$ which explains the bump in the density profile distribution seen in panel 2(a). A discontinuity in this location could mark the transition point between the outer atmosphere dominated by hydrogen and helium and deeper layers dominated by for example water or rocks. There are no other significant discontinuity regions, although the innermost region tends to have bigger slopes in general, and thus the discontinuity intensity also increases inward. In the same plot for Neptune (9(b)) the effects are reduced in visibility, but the same general trends exist with the peak being located slightly further out at $\sim 0.7R$. Note that here again, artefacts of the initial density profile generation method are visible, now as small cliffs around the inverses of multiples of 2 (Section 4 discusses this further).
@@
Modellers often assume a fixed number of density discontinuities (for example Nettelmann et al., 2013; Neuenschwander & Helled, 2022; Morf et al., 2024; Militzer, 2024). What number of such discontinuities should be chosen to account for all possible density profiles (given a certain criterion $c_{D}$)? To answer this question, we analyse the number of discontinuities compatible with a given $c_{D}$. Figure 10 shows the distribution of the total number of (minimally maximal and $c_{D}$-consistent) discontinuities in a given profile over all successful density profiles. We find that the number of discontinuities in one solution can be quite large, particularly in comparison to above mentioned published models who set the number of discontinuities to three. However, the results in Figure 10 clearly depend on the discontinuity criterion $c_{D}$. We hence investigate the influence of $c_{D}$ on the number of discontinuities in Figure 11. We show the average and upper 95th percentile number of discontinuities detected per successful density profile as a function of the criterion $c_{D}$. In addition, Table 3 lists all values from this figure numerically. Figure 11 and Table 3 should be read as follows: If we consider for example $c_{D}=$ 100 $\Delta\ell^{-1}$ kg m-3 to be a discontinuity, then the average number of discontinuities on Uranus is four, and 95% of all solutions have eight discontinuities or fewer. Alternatively, the statement ‘95% of Uranian density profiles have at most one discontinuity’ is only correct if we set $c_{D}\gtrsim$ 550 $\Delta\ell^{-1}$ kg m-3. Finally, we note that we consistently tend to find slightly more jumps for Neptune than Uranus, probably due to the larger uncertainty in Neptune’s gravity data, allowing for a larger parameter space of solutions.
@@ 4 Discussion
We begin by noting that our results depend on the measured values listed in Table 1. Some of these values remain uncertain, the rotation rates and the contributions of zonal winds for example remain unclear. Such fundamental uncertainties persist regardless of the used method. Below we focus on the limitations exclusive to our approach.
While our algorithm is able to efficiently and faithfully investigate the full space of planetary interior density profiles with an absolute minimum amount of assumptions, some caveats remain. Most importantly, our method for generating starting profiles has a higher variance (or degree of exploration) at some normalised radii compared to others. Listing 2 shows this generation method, the Binary Subdivision Algorithm.
Listing 2: The Binary Subdivision Algorithm
@@
The Binary Subdivision Algorithm starts by setting the density value at $r/R=1/2$. It does so by choosing uniformly at random from the entire range of possible values, subdividing the range into two halves in the process. It then recursively calls itself on those halves. As a result, some binary points in the range from $r/R=0$ to $r/R=1$ (such as 1/2, 1/4, 3/4, …) are decided earlier, and their bounds $\underline{y}$ and $\overline{y}$ for the (uniform) random generation are larger. Therefore, the variance at these points is higher than at others, and the generated starting density profiles tend to have kinks at these locations. These kinks can influence the final results, as can be seen in the small spikes that appear in Figure 5 and the cliffs in Figure 9. We nonetheless decided to use this method because it generates profiles sampled from the space of all possible density profiles such that the likelihood of generating a profile similar to another reference profile is equal everywhere. Appendix C.1 provides more details and a mathematically sound justification of the former statement. More practically speaking, we find that this limitation is acceptable because a large amount of starting profiles does not have these kinks and because the algorithm is able to remove these kinks after the initial generation. This is most apparent in Figure 9(a), where the algorithm has almost completely removed the cliffs for Uranus because of the tighter gravitational moment constraints that are incompatible with density discontinuities at inverse multiples of two in units of $r/R$. This can also be seen by directly comparing Figure 14 (Uranus), where there are fewer kinks or cliffs visible at inverse multiples of two, with Figure 15 (Neptune), where these kinks or cliffs are very prominent. We further provide a discussion on bias possibly introduced through the gradient method in Appendix C.2, which shows that our gradient descent method can be considered unbiased for the purpose of this work.
@@
While the linear approximation (see Appendix B) for the gradient was sufficient here, we have found that there are limitations in performance on harder problems. This is especially true if there are more measured gravitational moments or more precise measurements. There are several ways in which the algorithm could be modified accordingly. First, the optimisation scheme is always readily exchangeable. We used the Adam (Kingma & Ba, 2017) scheme, but any number of other descent schemes could be implemented. Due to the structure of the Theory of Figures, a full linear gradient was the most amenable choice, but depending on the method used to calculate gravitational moments, the problem could be approached with a stochastic gradient. This would alter the requirements, though Adam is still likely to perform very well. Second, the density profile parametrisation is also a viable candidate for modifications. Our parametrisation has the advantage of being simple to implement and easy to analyse for the gradient. However, it cannot structurally change the density profile in a meaningful way, the gradient cannot ’see’ directions of changes very well: Density profiles that are very similar to each other can have parameters ($p_{i}$ in Equation 3) that are very different. Different parametrisations could be preferred for this task, such as using orthogonal bases of polynomials for example.
@@
Finally, we note that a more comprehensive approach could eliminate the difficulties with the Theory of Figures and the need for a linear approximation entirely. Our algorithm was performance-bound exclusively by the implementation of the Theory of Figures we used, therefore a reduction in time usage there would have direct equal benefits. If a differentiable model could be trained to learn the Theory of Figures to within a certain acceptable loss, one would gain significant advantages in the accuracy of the gradients (and, presumably, the calculation of a single gradient would be quicker too). Alternatively, a non-differentiable machine model of the Theory of Figures would still have benefits in performance time, but would require zero-order methods. Additionally, an iterative approach where the average starting point moves towards the best performant point every so often could mean more time is spent resolving the space of acceptable solutions. However, this would not have served this work since our goal was the exploration of the entire space, not just the successful portion.
@@
Regarding future work, the analysis done on the solution space could be expanded further. One could include higher order or odd gravitational moments, especially in connection to dedicated Uranus and/or Neptune missions such as in Parisi et al. (2024); Mankovich et al. (2025). Chiefly, an analysis of the compositional and thermal structure of the planets based on the found density profiles could be performed. Special attention should be drawn to investigating the compositional structure around the hypothesised transition point between the hydrogen-helium rich outer layer and the below layer consisting of denser elements such as water or rocks. Morf & Helled (2025) do such a compositional and thermal analysis for profiles of Uranus and Neptune based on random starting profiles generated in a similar fashion to our models. However, their methods are limited by computational cost and hence yield only a handful of models (four for each planet). The key challenge would hence lie within ensuring high enough computational efficiency that does not come at the cost of physical self-consistency.
@@ 5 Summary and conclusions
We present a novel algorithm for efficiently and faithfully sampling planetary interior density profiles. Unlike conventional Markov Chain Monte Carlo (MCMC) approaches, our method employs a new optimisation-based strategy. We applied this algorithm to construct interior models of Uranus and Neptune, analysing the results from an extensive suite of simulations and comparing them to prior studies.
Our results are broadly consistent with previous studies, whose solutions typically fall within the 95% confidence interval of our model ensemble. However, these earlier models do not capture the full diversity of possible interior configurations, particularly in the deep interior. We also discuss limitations related to uneven sampling variance in our generation method and outline potential future directions, including refining the algorithm in various ways and expanding the model scope to include compositional analyses. Our key conclusions can be summarised as follows:
@@
    The median density (as a function of radius) for Uranus and Neptune shows gradual increases in density, consistent with composition gradients. The medians in particular maintain a consistent slope towards the centre and do not flatten off, with moderate central densities of $\sim$ 6 g cm-3 for Uranus and $\sim$ 7.5 g cm-3 for Neptune.
@@
    The solution space for Uranus and Neptune is least well constrained for the innermost region of the planets, with the 95% confidence interval reaching sizes of $\Delta\rho\sim$ 10 g cm-3 for Uranus and $\Delta\rho\sim$ 13 g cm-3 for Neptune.
@@
    Discontinuities with a slope bigger than 0.02 kg m-4 are generally rare.
@@
    Discontinuities with more moderate slopes (0.002 - 0.02 kg m-4) are frequent. The literature generally tends to only consider three or fewer discontinuities, which excludes a significant part of the solution space. If discontinuities are considered, we recommend adopting the values provided in Figure 11 or equivalently Table 3.
@@
    If at least one discontinuity is present, one is most likely to be found around $\sim 0.65R$ for Uranus and $\sim 0.7R$ for Neptune.

### v3

@@ 3.2 Derived quantities
Figure 6: Pressure profiles. Weighted distribution of pressure profiles derived from all successful density profiles. The inset shows a zoomed-in view of the outermost region, from $r/R=1$ to $r/R=0.7$. Note that the zoom is not aspect-preserving. The colour scale shows the relative frequency over all distributions of a certain pressure value at a given radius. The coloured lines are as before in Figures 2(a) and 3(a).
@@
In this Section, we present quantities derived from the successful density profiles via the ToF. Figure 6 shows the distribution of pressure profiles with a logarithmic frequency scale. Figure 7 shows the distribution of the normalised moments of inertia, and Figure 8 shows the distribution of the flattening ratios.
We find that the pressure profiles in Figure 6 agree well with previous studies and are more constrained for Uranus compared to Neptune. The pressure profiles were calculated from the density profiles under the assumption of hydrostatic equilibrium, $\nabla p=\rho\nabla U$, where $U$ is the total (gravitational and rotational) potential of the planet.
The normalised moments of inertia depicted in Figure 7 are defined as $\lambda=I/(mR_{\text{mean}}^{2})$, where $I$ is the axial moment of inertia,
$I=\int_{V}\rho r^{2}_{\perp}\mathrm{d}V,$ (7)
and $r_{\perp}$ denotes the perpendicular radial component with respect to the given axis, usually the axis of rotation. Further, $R_{\text{mean}}$ is equal to $l_{\text{max}}$, the outermost spheroid volumetric mean radius (see Section 2). We find that our average value for Uranus of 0.22595 agrees well with previous findings from Neuenschwander & Helled (2022), who found a range of 0.22594–0.22670. Our average value for Neptune of 0.23597 is lower than their finding of 0.23727–0.23900 for the standard model, though our more frequent values agree with the lower end of their range.
@@
To compare our findings, we calculated the equatorial-to-mean volumetric radii ratios based on the values from Seidelmann et al. (2007). They are $(25559\pm 4)/(25362\pm 7)$ for Uranus and $(24764\pm 15)/(24622\pm 19)$ for Neptune. The resulting flattening ratios are $r_{f}=1.0078\pm 0.0003$ and $r_{f}=1.0058\pm 0.0010$ for Uranus and Neptune, respectively. Our inferred average value of 1.0068 in Figure 8 for Uranus is lower compared to Seidelmann et al. (2007), while our average of 1.0063 for Neptune is within their uncertainties. Our ToF implementation is benchmarked against Wisdom & Hubbard (2016), which shows that our flattening ratios should be accurate up to a relative error of $10^{-3}-10^{-4}$. This is significant given the uncertainties from Seidelmann et al. (2007) for Uranus and can therefore explain this deviation. Furthermore, the uncertainties given by Seidelmann et al. (2007) are quite conservative, especially given the uncertain rotation rates of these planets (for example Helled et al., 2010).
@@ 3.3 Discontinuities
Figure 9: Weighted discontinuity ‘intensity’ for Uranus (9(a)) and Neptune (9(b)). The figures depict the average height, $\overline{\Delta\rho}$, divided by the average width, $\overline{\Delta r}$, of the discontinuities. For example, Uranus discontinuities with an average location of $r/R=0.5$ possess $\overline{\Delta\rho/\Delta r}\sim 0.5\Delta\ell^{-1}$ kg m-3. Profiles with no discontinuities also contribute to the averages with a value zero.
In this section, we present an analysis on discontinuities of successful density profiles. First, a ‘discontinuity criterion’ ($c_{D}$) must be chosen in order to distinguish discontinuities from regular slopes. Unless otherwise specified, all results were obtained with $c_{D}=$ 100 $\Delta\ell^{-1}$ kg m-3. As a reminder, $\Delta\ell$ is equal to 24960 m for Uranus and 24186 m for Neptune (rounded to the nearest meter). Therefore, $c_{D}$ has a value of 4.01 $\cdot$ 10-3 kg m-4 for Uranus and 4.13 $\cdot$ 10-5 kg m-4 for Neptune (rounded to two significant figures). As a second distinguishing feature, we only considered discontinuities that are ‘minimally maximal’. A precise definition of the term minimally maximal is given in Appendix D. In a nutshell, minimally maximal specifies that the (radial) width of a discontinuity can neither be extended (because it would add sections of the density profile below $c_{D}$) nor shortened (because it would remove sections of the density profile exceeding $c_{D}$).
Figure 9 shows the average discontinuity ‘intensity’ as a function of radius for all successful density profiles for both Uranus and Neptune. More precisely, it shows the average discontinuity height, $\overline{\Delta\rho}$, divided by the average discontinuity width, $\overline{\Delta r}$, for discontinuities that are minimally maximal and in agreement with $c_{D}$ at their respective average normalised radius, $r/R$. For example, Figure 9(a) shows that minimally maximal and $c_{D}$-consistent discontinuities in Uranus with an average location of $r/R=0.5$ possess an average slope of $\overline{\Delta\rho/\Delta r}\sim 0.5\Delta\ell^{-1}$ kg m-3. We note that profiles that have no minimally maximal and $c_{D}$-consistent discontinuity at any given location also contribute to the averages with a discontinuity intensity value of zero. Separately, we show $\overline{\Delta\rho}$ and $\overline{\Delta r}$ at their respective average normalised radii, $r/R$, in Figures 14 for Uranus and 15 for Neptune.
@@
For Uranus (Figure 9(a)), we observed a discontinuity intensity peak at $\sim 0.65R$, which explains the bump in the density profile distribution seen in panel 2(a). A discontinuity in this location could mark the transition point between the outer atmosphere dominated by hydrogen and helium and deeper layers dominated by, for example, water or rocks. There are no other significant discontinuity regions, although the innermost region tends to have bigger slopes in general, and thus the discontinuity intensity also increases inward. In the same plot for Neptune (9(b)) the effects are reduced in visibility, but the same general trends exist, with the peak being located slightly further out at $\sim 0.7R$. We note that here as well artefacts of the initial density profile generation method are visible, but this time as small cliffs around the inverses of multiples of two (we discuss this further in Section 4).
@@
Modellers often assume a fixed number of density discontinuities (for example Nettelmann et al., 2013; Neuenschwander & Helled, 2022; Morf et al., 2024; Militzer, 2024). However, this leads to the question of how many of these discontinuities should be chosen to account for all possible density profiles (given a certain criterion $c_{D}$). To answer this, we analysed the number of discontinuities compatible with a given $c_{D}$. Figure 10 shows the distribution of the total number of (minimally maximal and $c_{D}$-consistent) discontinuities in a given profile over all successful density profiles. We find that the number of discontinuities in one solution can be quite large, particularly in comparison to the above-mentioned published models, which set the number of discontinuities to three. However, the results in Figure 10 clearly depend on the discontinuity criterion, $c_{D}$. We hence investigated the influence of $c_{D}$ on the number of discontinuities in Figure 11. We show the average and upper 95th percentile number of discontinuities detected per successful density profile as a function of the criterion $c_{D}$. In addition, Table 3 lists all values from this figure numerically. Figure 11 and Table 3 should be read as follows: If we consider, for example, $c_{D}=$ 100 $\Delta\ell^{-1}$ kg m-3 to be a discontinuity, then the average number of discontinuities on Uranus is four, and 95% of all solutions have eight discontinuities or fewer. Alternatively, the statement ‘95% of Uranian density profiles have at most one discontinuity’ is only correct if we set $c_{D}\gtrsim$ 550 $\Delta\ell^{-1}$ kg m-3. Finally, we note that we consistently tend to find slightly more jumps for Neptune than Uranus, probably due to the larger uncertainty in Neptune’s gravity data, which allows for a larger parameter space of solutions.
@@ 4 Discussion
We begin by noting that our results depend on the measured values listed in Table 1. Some of these values remain uncertain. The rotation rates and the contributions of zonal winds, for example, remain unclear. Such fundamental uncertainties persist regardless of the method used. Below we focus on the limitations exclusive to our approach.
While our algorithm is able to efficiently and faithfully investigate the full space of planetary interior density profiles with an absolute minimum amount of assumptions, some caveats remain. Most importantly, our method for generating starting profiles has a higher variance (or degree of exploration) at some normalised radii compared to others. Listing 2 shows our generation method: the binary subdivision algorithm. The binary subdivision algorithm starts by setting the density value at $r/R=1/2$. It does so by uniformly choosing at random from the entire range of possible values and subdividing the range into two halves in the process. It then recursively calls itself on those halves.
Listing 2: The binary subdivision algorithm.
@@
As a result, some binary points in the range from $r/R=0$ to $r/R=1$ (for example 1/2, 1/4, 3/4) are decided earlier, and their bounds, $\underline{y}$ and $\overline{y}$, for the (uniform) random generation are larger. Therefore, the variance at these points is higher than at others, and the generated starting density profiles tend to have kinks at these locations. These kinks can influence the final results, as can be seen in the small spikes that appear in Figure 5 and the cliffs in Figure 9. We nonetheless decided to use this method because it generates profiles sampled from the space of all possible density profiles such that the likelihood of generating a profile similar to another reference profile is equal everywhere. Appendix C.1 provides more details and a mathematically sound justification of the former statement. More practically speaking, we find that this limitation is acceptable because many of the starting profiles do not have these kinks and because the algorithm is able to remove the kinks after the initial generation. This is most apparent in Figure 9(a), where the algorithm has almost completely removed the cliffs for Uranus because of the tighter gravitational moment constraints that are incompatible with density discontinuities at the inverse multiples of two in units of $r/R$. This can also be seen by directly comparing Figure 14 (Uranus), where there are fewer kinks or cliffs visible at inverse multiples of two, with Figure 15 (Neptune), where these kinks or cliffs are very prominent. We provide a discussion on the bias possibly introduced through the gradient method in Appendix C.2, which shows that our gradient descent method can be considered unbiased for the purpose of this work.
@@
While the linear approximation (see Appendix B) for the gradient was sufficient here, we find that there are limitations in performance on harder problems. This is especially true if there are more measured gravitational moments or more precise measurements. There are several ways in which the algorithm could be modified accordingly. First, the optimisation scheme is always readily exchangeable. We used the Adam (Kingma & Ba, 2017) scheme, but any number of other descent schemes could be implemented. Due to the structure of the ToF, a full linear gradient was the most amenable choice, but depending on the method used to calculate gravitational moments, the problem could be approached with a stochastic gradient. This would alter the requirements, though Adam is still likely to perform very well. Second, the density profile parametrisation is also a viable candidate for modifications. Our parametrisation has the advantage of being simple to implement and easy to analyse for the gradient. However, it cannot structurally change the density profile in a meaningful way, as the gradient cannot ’see’ directions of changes very well. Density profiles that are very similar to each other can have parameters ($p_{i}$ in Equation 3) that are very different. Different parametrisations could be preferred for this task, such as using orthogonal bases of polynomials, for example.
@@
Finally, we note that a more comprehensive approach could eliminate the difficulties with the ToF and the need for a linear approximation entirely. Our algorithm was performance-bound exclusively by the implementation of the ToF we used; therefore, a reduction in time usage there would have direct equal benefits. If a differentiable model could be trained to learn the ToF to within a certain acceptable loss, one would gain significant advantages in the accuracy of the gradients (and, presumably, the calculation of a single gradient would be quicker too). Alternatively, a non-differentiable machine model of the ToF would still have benefits in performance time but would require zero-order methods. Additionally, an iterative approach where the average starting point moves towards the best performant point every so often could mean more time is spent resolving the space of acceptable solutions. However, this would not have served this work since our goal was the exploration of the entire space, not just the successful portion.
@@
Regarding future work, the analysis done on the solution space could be expanded further. One could include higher-order or odd-gravitational moments, especially in connection to dedicated Uranus and/or Neptune missions such as in Parisi et al. (2024); Mankovich et al. (2025). Chiefly, an analysis of the compositional and thermal structure of the planets based on the found density profiles could be performed. Special attention should be given to investigating the compositional structure around the hypothesised transition point between the hydrogen-helium rich outer layer and the below layer consisting of denser elements such as water or rocks. Morf & Helled (2025) perform such a compositional and thermal analysis for profiles of Uranus and Neptune based on random starting profiles generated in a fashion similar to our models. However, their methods are limited by computational cost and hence yield only a handful of models (four for each planet). Hence, the key challenge would lie within ensuring a high enough computational efficiency that does not come at the cost of physical self-consistency.
@@ 5 Summary and conclusions
We have presented a novel algorithm for efficiently and faithfully sampling planetary interior density profiles. Unlike conventional MCMC approaches, our method employs a new optimisation-based strategy. We used this algorithm to construct interior models of Uranus and Neptune, and we analysed the results from an extensive suite of simulations and compared them to prior studies.
Our results are broadly consistent with previous studies, whose solutions typically fall within the 95% confidence interval of our model ensemble. However, the earlier models do not capture the full diversity of possible interior configurations, particularly in the deep interior. We also discussed the limitations related to uneven sampling variance in our generation method and outlined potential future directions, including refining the algorithm in various ways and expanding the model scope to include compositional analyses. Our key conclusions are summarised as follows:
@@
    The median density (as a function of radius) for Uranus and Neptune shows gradual increases in density consistent with composition gradients. In particular, the medians maintain a consistent slope towards the centre and do not flatten off, with moderate central densities of $\sim$ 6 g cm-3 for Uranus and $\sim$ 7.5 g cm-3 for Neptune.
@@
    The solution space for Uranus and Neptune is least well constrained for the innermost region of the planets, with the 95% confidence interval reaching sizes of $\Delta\rho\sim$ 11 g cm-3 for Uranus and $\Delta\rho\sim$ 13 g cm-3 for Neptune.
@@
    Discontinuities with a slope greater than 0.02 kg m-4 are generally rare.
@@
    Discontinuities with more moderate slopes (0.002 - 0.02 kg m-4) are frequent. The literature generally tends to only consider three or fewer discontinuities, which excludes a significant part of the solution space. If discontinuities are considered, we recommend adopting the values provided in Figure 11 or equivalently those in Table 3.
@@
    If at least one discontinuity is present, it is most likely to be found around $\sim 0.65R$ for Uranus and $\sim 0.7R$ for Neptune.